solve using addition and multiplication priinciple
7-10z+4>-11z+5
please note, there's a line under the > symbol. I do not know how to put a line under the > sumbol

Answer by mathmandotcom(31)

7-10z+4>=-11z+5
11-10z>=-11z+5 subtract 5
6-10z>=-11z; add 10z
6>=-1z; mult by -1 and reverse the symbol
-6<=z or z>=-6